Chapter 2, Section 2.4, Question E62 Unstandardizing. Find the value of x that was converted to the given z-score. a. z 3, mean 20, standard deviation 8 b. z-1, mean 46, standard deviation 5 c. z- -3.8, mean 200, standard deviation 10 d, z-2.5, mean-10, standard deviation 0.8
Answer #1

z=(x-µ)/σ
so, X=zσ+µ=
a) = X=3*8+20 = 44

b) X=zσ+µ

x= -1*5+46 = 41

c) X=zσ+µ

x=-3.8*10+200=162

d) x = 2.5*0.8-10=-8